find the angle made by the vector 4i+3j+5k on x axis​
1
Expert's answer
2020-03-09T11:16:06-0400

Given a vector "{\bf A}=4\hat i+3\hat j+5\hat k." The angle between "{\bf A}" and positive "x" axis is given by


"\theta=\cos^{-1}\frac{A_x}{A}=\cos^{-1}\frac{A_x}{\sqrt{A_x^2+A_y^2+A_z^2}}""\theta=\cos^{-1}\frac{4}{\sqrt{4^2+3^2+5^2}}=\cos^{-1}\frac{4}{5\sqrt{2}}=0.566=55.6^{\circ}"